OpenStudy (anonymous):

WILL BECOME FAN AND AWARD MEDAL!!! Make up any two functions f(x) and g(x), defined in terms of the variable x. For example, you might define f(x) = x – 15 and g(x) = 3x2. (Use different examples.) Then provide the compositions of the two functions f(g(x)) and g(f(x)). Which composite function has the greatest value for x = 10? Remember no right answer exists. Your answer completely depends on the functions you select.

OpenStudy (anonymous):

did you pick two functions?

OpenStudy (anonymous):

go ahead and make two up they can be anything, but you don't want to make them too hard, so stick with something simple

OpenStudy (anonymous):

once you have them we can do this easy

OpenStudy (anonymous):

f(x)= x+7 and g(x)= 1/x+7 were the functions I thought to use for this.

OpenStudy (anonymous):

\[f(x)=x+7,g(x)=\frac{1}{x+7}\] like that?

OpenStudy (anonymous):

or was \(g(x)=\frac{1}{x}+7\) either way we can do it

OpenStudy (anonymous):

you pick

OpenStudy (anonymous):

the second one was how I meant to put the second function. sorry ^.^"

OpenStudy (anonymous):

ok brave to use a denominator but no problem lets compute \[f(g(x))\] first

OpenStudy (anonymous):

always work from the inside out \[f(g(x))=f(\frac{1}{x}+7)\] is a fist step

OpenStudy (anonymous):

then since \(f9\heartsuit)=\heartsuit +7\) you have ' \[f(\frac{1}{x}+7)=\frac{1}{x}+7+7\] or more simply \[\frac{1}{x}+14\]

OpenStudy (anonymous):

\[f(x)=x+7\] and \[g(x)= \frac{ 1}{ x } +7\] this is how they should look

OpenStudy (anonymous):

yeah got it

OpenStudy (anonymous):

that makes \[f(g(x))=f(\frac{1}{x}+7)=\frac{1}{x}+7+7=\frac{1}{x}+14\]

OpenStudy (anonymous):

Sorry. My computer is messing up so I'm having a hard time seeing your responses. I'm not sure why open study isn't working properly for me, but it isn't...

OpenStudy (anonymous):

yeah for me too

OpenStudy (anonymous):

did you see the last one?

OpenStudy (anonymous):

Yes! And thank you for your help! I really do appreciate it :3

OpenStudy (anonymous):

for the next one it is a bit tricker

OpenStudy (anonymous):

\[g(f(x))=g(x+7)=\frac{1}{x+7}+7\]

OpenStudy (anonymous):

you can replace \(x\) by 10 in both of those for \[g(f(10))\] you will get \[\frac{1}{10+7}=7=\frac{1}{17}+7=7\tfrac{1}{17}\]

OpenStudy (anonymous):

you mean to get the result for Which composite function has the greatest value for x=10?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

Sorry >.< late response again. But I believe I understand the material. ^w^

OpenStudy (anonymous):

ok good stay cool

OpenStudy (anonymous):

You too! ^w^ Thanks for the help @satellite73 :3